OpenStudy (anonymous):

Determine the slope of the line represented by the equation: y = 1/4 x + 10

OpenStudy (anonymous):

the multiplier of x is the slope, 1/4. hope helps

OpenStudy (anonymous):

\(y=mx+b\) where \(m\) is the slope

OpenStudy (anonymous):

what about when its written ... -10x+1y=-1

OpenStudy (anonymous):

you turn it into y=mx+b form...so basically just make it equal to y

OpenStudy (anonymous):

Hint: start by adding 10x to both sides

OpenStudy (anonymous):

\[ -10x+1y=-1\]\[y=10x+1\]

OpenStudy (anonymous):

and 10 would be the slope?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

no problem!